OpenStudy (anonymous):

Find the geometric sequence with three means between 1/6 and 1/486

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (anonymous):

Also this is my first time using this site, so sorry if I do something wrong

OpenStudy (watchmath):

you mean with three terms?

OpenStudy (anonymous):

Yes

OpenStudy (watchmath):

1/6, 1/54, 1/486 and the ratio is 1/9

OpenStudy (anonymous):

No its 5 terms in all, you have to find 3 of them

OpenStudy (anonymous):

So like 1/6,___,____,____,1/486

OpenStudy (anonymous):

Anybody???

OpenStudy (watchmath):

let say the ratio is r (which we will determined later). Then to get 1/486, we need to multiply 1/6 with r four times. Or (1/6)r^2=1/486. Now try to solve for r

OpenStudy (watchmath):

sorry \(\frac{1}{6}r^4=\frac{1}{486}\)

OpenStudy (anonymous):

So r=about 1/3?

OpenStudy (anonymous):

Or am I lost?

OpenStudy (watchmath):

yes, excellent! Now how to use that information to get your sequence?

OpenStudy (anonymous):

You sub in 1/3 for r?

OpenStudy (anonymous):

Got it! Thanks for the help :)